The total revenue R is directly proportional to the numver of units sold x. When 25 units are sold, the revenue is Php 30,000. Find the revenue when 58 units were sold?
Answer by fractalier(6550)

Direct proportion means
R = kx.
If R = 30000 when x = 25, we can find k...
k = R/x = 30000/25 = 1200
Thus we have
R = 1200x
If x = 58 we have
R = 1200(58) = 69600